# SAKE: Towards Editing Auditory Attribute Knowledge of Large Audio-Language Models

### v1

@@ 4 Experimental Settings
We experiment on two LALMs, DeSTA2.5-Audio (Lu et al., 2025b) and Qwen2-Audio-Instruct (Chu et al., 2024), chosen for their strong benchmark performance (Huang et al., 2025; Yang et al., 2025b; Sakshi et al., 2025; Chen et al., 2024b; Lu et al., 2025c). We apply greedy decoding and assess editing methods under two settings: single editing and sequential editing.
Figure 2: Example of sequential editing. For comparability, only the first five edits are evaluated, with gaps capped at five. For example, evaluating Edit 2 immediately results in a gap of 0, whereas after Edit 10, only Edit 5 can be evaluated, with a gap of 5 under this rule.
@@ 4.1 Editing Methods
We evaluate seven editing methods, focusing on their effectiveness in modifying abstract auditory attribute knowledge. Below, we briefly introduce these methods, with the details provided in Appendix C.
Fine-tuning is a common approach for adapting pre-trained models to new knowledge. Following prior work on LVLMs (Cheng et al., 2023; Huang et al., 2024; Zhang et al., 2024), we compare fine-tuning two parts in LALMs: the last layer of the LLM backbone, denoted as FT (LLM), and the modality connector between the audio encoder and the LLM backbone, denoted as FT (Audio). Knowledge Editor (KE) (De Cao et al., 2021) and MEND (Mitchell et al., 2022) trains a hypernetwork to transform the fine-tuning gradients into parameter updates for the edit. UnKE (Deng et al., 2025) optimizes specific neurons of the chosen layers to produce the edit target. In-Context Knowledge Editing (IKE) (Zheng et al., 2023) leverages in-context learning (ICL) to enforce knowledge updates. We consider two variants: Instruction-based IKE (I-IKE), which encodes edits solely through natural language instructions in system prompts, and Instruction+Example IKE (IE-IKE), which provides auditory examples retrieved from the training set as dialog context.
@@ 4.2 Single Editing and Sequential Editing
Single editing evaluates the performance of editing a single piece of knowledge, whereas sequential editing evaluates the performance after applying a sequence of edits continuously on different knowledge, which better reflects real-world scenarios.
For sequential editing, we construct ten independent sequences, each with ten editing instances. An editing sequence is denoted as $\mathcal{S}=\{(a_{e}^{(t)},x_{e}^{(t)},y_{o}^{(t)},y_{e}^{(t)})\}_{t=1}^{10}$, where $(a_{e}^{(t)},x_{e}^{(t)},y_{e}^{(t)})$ is sampled from $\mathcal{D}_{edit}$ and the corresponding original label $y_{o}^{(t)}$ is retrieved from our audio dataset. Here, $t$ indexes the order of edits within the sequence.
@@ 4.3 Evaluator
Because LALMs often generate descriptive responses, we adopt LLM-as-a-judge (Chiang & Lee, 2023) to compute the metrics introduced in Sec. 3.2. In particular, we employ GPT-5 mini (gpt-5-mini-2025-08-07) as the evaluator, focusing on correctness and consistency.
For the reliability, generality, and portability metrics, we assess whether the edited model’s responses correctly align with the ground truth. For locality, we examine whether the edited model preserves consistency with the original model’s outputs. To further validate the quality of the LLM-based judgments, we conduct human evaluation on 420 randomly selected samples. The results show overall 98.10% agreement with the LLM evaluator, demonstrating its robustness. Additional details and the evaluation prompts are provided in Appendix D.
Table 1: The four metrics (%) of the editing methods on the two models. Avg. indicates the average performance across all types of the corresponding metric. The best and second-best results are highlighted in bold and underlined, respectively.
Method Reliability Generality Audio Locality Text Locality Portability
Avg. Type 1 Type 2 Type 3 Avg. Type 1 Type 2 Type 3 Type 4
DeSTA2.5-Audio
FT (LLM) 99.75 98.75 99.67 99.00 97.58 65.11 88.08 15.56 74.67 69.75 82.58 19.42
FT (Audio) 99.50 86.14 96.75 84.92 76.75 68.09 78.67 48.11 76.17 64.42 100.00 55.50
KE 99.58 99.17 99.25 99.25 99.00 79.29 96.33 43.89 76.33 91.75 92.50 18.42
MEND 95.33 95.00 95.83 95.17 94.00 71.47 93.42 17.22 74.67 87.00 92.08 19.42
UnKE 96.30 89.28 89.67 94.25 83.92 56.04 67.67 10.56 71.92 62.67 87.38 18.73
I-IKE 73.00 61.47 64.67 59.58 60.17 65.40 79.42 50.67 73.67 54.17 62.08 71.42
IE-IKE 40.58 39.61 41.50 38.67 38.67 58.76 70.08 49.89 70.75 42.08 56.25 34.58
Qwen2-Audio
FT (LLM) 100.00 99.94 100.00 100.00 99.83 67.42 91.67 10.44 83.33 70.00 74.58 24.67
FT (Audio) 100.00 81.86 99.00 77.17 69.42 90.53 96.83 80.11 90.00 92.58 100.00 50.67
KE 95.50 86.67 92.00 87.67 80.33 83.47 89.83 61.44 87.25 89.83 84.92 27.58
MEND 100.00 95.33 98.92 95.92 91.17 83.27 98.50 49.33 85.42 91.33 86.75 27.17
UnKE 98.60 98.53 99.08 98.83 97.67 67.49 91.42 12.34 82.92 69.50 71.58 30.18
I-IKE 10.33 7.11 10.33 5.67 5.33 87.51 94.75 89.33 93.00 73.42 55.00 28.92
IE-IKE 8.00 6.58 8.50 5.75 5.50 82.89 91.25 86.56 89.67 65.00 50.00 27.50
@@ 5.1 Single Editing
The main results of different knowledge editing methods on SAKE are shown in Table 1. Detailed results for each auditory attribute and relevant discussions are provided in Appendix E.
Reliability.
@@
For both LALMs, most editing methods achieve high reliability, with FT (LLM), i.e., fine-tuning the LLM backbone, consistently yielding the best scores. In contrast, I-IKE and IE-IKE perform poorly, reaching only 73.00% and 40.58% on DeSTA2.5-Audio, and 10.32% and 8.00% on Qwen2-Audio. Interestingly, I-IKE outperforms IE-IKE on both models despite the latter using additional auditory examples. We attribute this to LALMs’ limited in-context learning ability: they struggle to handle multi-audio inputs and leverage examples, unlike in LLMs (Zheng et al., 2023) and LVLMs (Cheng et al., 2023; Huang et al., 2024) where such methods are effective.
Generality.
Compared with reliability, editing methods yield lower average generality scores. KE performs best on DeSTA2.5-Audio, while FT (LLM) leads on Qwen2-Audio. Baselines generally perform well on type 1 (textual neighborhood) but decline on type 2, which requires generalizing to similar but not identical auditory inputs with the same labels. Type 3, combining both text and auditory neighborhoods, proves most challenging, showing that current methods struggle to extend edited knowledge to auditory modalities as effectively as to text.
Interestingly, FT (LLM) consistently outperforms FT (Audio) on generality despite their comparable reliability scores. This suggests that while fine-tuning either the LLM backbone or the modality connector ensures success on the editing instances themselves, the ability to generalize to similar inputs differs substantially, with training on the LLM backbone offering better generalizability.
Audio Locality.
Most methods perform poorly, with only limited success: KE preserves knowledge best on DeSTA2.5-Audio, while FT (Audio) performs better on Qwen2-Audio. This shows that maintaining stability in the auditory domain remains challenging despite advances in model editing.
Among audio locality types, type 2 locality, where knowledge within the same attribute but unrelated to the edit is examined, is hardest to preserve, more so than knowledge of other attributes (type 1) or the edit target itself (type 3). This suggests that edits on one aspect of an attribute may inadvertently affect other aspects of the same attribute, reflecting a degree of intra-attribute entanglement of auditory attribute knowledge. Notably, FT (LLM) performs much worse than FT (Audio) on type 2, indicating that despite better generalization, it fails to retain unrelated knowledge within the same attribute. By contrast, knowledge across different attributes (i.e., type 1) is less affected, indicating that cross-attribute boundaries tend to be more stable to preserve during editing.
Focusing on type 4, which evaluates the preservation of general auditory processing capability, most editing methods struggle to retain LALMs’ general abilities, suggesting a trade-off between effective editing and maintaining broad competence. KE and MEND perform best, likely due to their regularization that emphasizes locality in hypernetwork training.
Although I-IKE and IE-IKE do not achieve high reliability and generality, they still preserve audio locality to some extent. This can be attributed to the limited in-context learning ability of DeSTA2.5-Audio and Qwen2-Audio, which reduces the extent to which editing propagates. However, since DeSTA2.5-Audio possesses stronger instruction-following capability (Lu et al., 2025b), it is more susceptible to perturbations during editing, leading to lower locality scores than Qwen2-Audio.
Text Locality.
FT (Audio) achieves 100% since it fine-tunes only the modality connector, which is triggered solely by auditory inputs and leaves text-only capabilities intact. FT (LLM) tunes the LLM backbone, impairing text-only tasks and yielding much lower scores. KE and MEND preserve text locality better through regularization on irrelevant samples during hypernetwork training. UnKE disrupts text-based knowledge, especially on Qwen2-Audio, while IKE variants also struggle, likely due to sensitivity to contextual information in text-only scenarios, even when irrelevant.
Portability.
Overall, current editing methods do not guarantee portability when modifying auditory attribute knowledge. Parameter-updating approaches generally overlook this aspect, which poses a major challenge for real-world deployment, where extensive related knowledge may need to be updated simultaneously. Among these methods, FT (Audio) achieves the most balanced performance on both LALMs, suggesting that training the modality connector can better integrate edited auditory knowledge with the internal world knowledge encoded in the LLM backbone.
Interestingly, I-IKE performs best on DeSTA2.5-Audio. This may be because portability is closely tied to reasoning, and the IKE variants benefit when the underlying model has sufficient reasoning proficiency. Since DeSTA2.5-Audio demonstrates substantially stronger reasoning ability than Qwen2-Audio (Lu et al., 2025b), the IKE variants perform better on it. However, due to DeSTA2.5-Audio’s still-limited multi-audio processing ability, IE-IKE remains less effective than I-IKE.
Summary.
In sum, while most editing methods can adjust LALMs to produce the desired knowledge on edited instances, they struggle to generalize these changes to equivalent cases, especially for auditory inputs. They also struggle to extend edits consistently to interconnected knowledge. In addition, they often fail to preserve original knowledge within the same attribute, which highlights the difficulty of preventing unintended perturbations. Among existing methods, fine-tuning, especially FT (Audio), remains a strong baseline across our evaluations. Editing auditory attribute knowledge, therefore, remains highly challenging and calls for methods tailored to auditory modalities.
((a)) Sequential editing results of DeSTA2.5-Audio by edit gap (0–5).
((b)) Sequential editing results of Qwen2-Audio by edit gap (0–5).
@@ 5.2 Sequential Editing
Figure 3(a) and 3(b) show sequential editing results on DeSTA2.5-Audio and Qwen2-Audio.
Reliability and Generality. Most methods decline in reliability and generality as the gap increases, indicating a tendency to forget previously edited auditory knowledge. On DeSTA2.5-Audio, MEND and KE deteriorate rapidly after a few edits, often degenerating into repetitive outputs, with examples provided in Appendix G. In contrast, although weaker in single-editing settings, the IKE variants remain comparatively stable across larger gaps, yielding stronger long-term reliability and generality on DeSTA2.5-Audio, particularly for I-IKE.
Locality. Most methods do not show the sharp declines seen in reliability and generality, suggesting sequential setups have a weaker impact on knowledge preservation. However, as discussed in Sec. 5.1, their robustness still leaves considerable room for improvement, as none of the methods fully address the challenge of consistent knowledge preservation. Notable exceptions are KE and MEND on DeSTA2.5-Audio and KE and IE-IKE on Qwen2-Audio, which show marked declines due to degeneration issues and the models’ limited in-context learning ability.
Portability. Similarly, the portability metric does not show significant declines under sequential editing. I-IKE and FT (Audio) consistently outperform other methods on DeSTA2.5-Audio and Qwen2-Audio, respectively, in line with the single-edit results in Table 1. This suggests that their ability to update interconnected knowledge remains relatively stable across several edits. Nevertheless, the overall performance of all methods remains unsatisfactory, underscoring the need for further improvement.

### v2

@@ 4 Experimental Settings
We use three LALMs, DeSTA2.5-Audio (Lu et al., 2025b), Qwen2-Audio-Instruct (Chu et al., 2024), and Audio Flamingo 3 (Ghosh et al., 2025a), chosen for their strong benchmark performance (Huang et al., 2025; Yang et al., 2025b; Sakshi et al., 2025; Chen et al., 2024b; Lu et al., 2025c). We refer to them as DeSTA, Qwen, and AF, respectively. Details about the model choice are in Appendix C. We apply greedy decoding and assess editing methods under two settings: single editing and sequential editing.
@@ 4.1 Editing Methods
We evaluate eight knowledge editing methods that are widely adopted in prior works (Cheng et al., 2023; Huang et al., 2024; Zhang et al., 2025; Du. et al., 2025; Chen et al., 2025; Ma et al., 2025a), focusing on their effectiveness in modifying abstract auditory attribute knowledge. These methods span diverse methodological paradigms, including fine-tuning-based, hypernetwork-driven, optimization-centric, memory-augmented, and in-context editing approaches. This coverage enables systematic evaluation across a wide range of editing methods. We provide an overview of these methods below, with implementation details deferred to Appendix 7.
Fine-tuning is a common approach for adapting pre-trained models to new knowledge. Following prior work on LVLMs (Cheng et al., 2023; Huang et al., 2024; Zhang et al., 2025), we compare fine-tuning two parts in LALMs: the last layer of the LLM backbone (FT (LLM)) and the modality connector between the audio encoder and the LLM backbone (FT (Audio)). Knowledge Editor (KE) (De Cao et al., 2021) and MEND (Mitchell et al., 2022) trains a hypernetwork to transform the fine-tuning gradients into parameter updates for the edit. UnKE (Deng et al., 2025) optimizes specific neurons of the chosen layers to produce the edit target. In-Context Knowledge Editing (IKE) (Zheng et al., 2023) leverages in-context learning (ICL) to enforce knowledge updates. We consider two variants: Instruction-based IKE (I-IKE), which encodes edits solely through natural language instructions in system prompts, and Instruction+Example IKE (IE-IKE), which provides auditory examples retrieved from the training set. WISE (Wang et al., 2024b) edits models using a dual-memory design that keeps pretrained FFN weights intact while storing edit-specific updates in a lightweight side memory.
@@ 4.2 Single Editing and Sequential Editing
Single editing evaluates the performance of editing a single piece of knowledge; sequential editing evaluates the performance after applying several edits continuously on different knowledge, which better reflects real-world scenarios.
For sequential editing, we construct ten independent sequences, each with ten editing instances sampled from the data used for single editing. An editing sequence is denoted as $\mathcal{S}=\{(a_{e}^{(t)},x_{e}^{(t)},y_{o}^{(t)},y_{e}^{(t)})\}_{t=1}^{10}$, where $(a_{e}^{(t)},x_{e}^{(t)},y_{e}^{(t)})$ is sampled from $\mathcal{D}_{edit}$ and the original label $y_{o}^{(t)}$ is retrieved from our audio dataset. Here, $t$ indexes the order of edits within the sequence.
@@
Figure 2: Example of sequential editing. For comparability, only the first five edits are evaluated, with the evaluation gap capped at five. For instance, Edit 2 is evaluated with a gap of 0, while after Edit 10, only Edit 5 is evaluated with a gap of 5.
@@ 4.3 Evaluator
Because LALMs often generate descriptive responses, we adopt LLM-as-a-judge (Chiang and Lee, 2023) to compute the metrics introduced in Sec. 3.2, using GPT-5 mini as the evaluator. To validate these judgments, we conduct human evaluation on 420 randomly sampled cases, achieving 98.10% agreement, indicating strong robustness. Additional details and prompts are provided in Appendix E.
Table 2: Performance (%) of editing methods on three models. Generality and audio locality scores are averaged across all attributes and evaluation types. Best and second-best results on individual metrics are shown in bold and underlined, respectively.
Model Metrics FT (LLM) FT (Audio) KE MEND UnKE I-IKE IE-IKE WISE
DeSTA Reliability ($\uparrow$) 99.75 99.50 99.58 95.25 96.09 64.25 40.58 100.00
Generality ($\uparrow$) 98.75 86.06 99.17 94.94 87.42 61.08 39.61 100.00
Audio Locality ($\uparrow$) 64.93 68.13 79.49 71.44 57.18 65.49 58.89 38.90
Text Locality ($\uparrow$) 82.67 100.00 93.08 92.00 88.50 61.92 56.67 83.08
Portability ($\uparrow$) 18.83 55.33 17.33 18.08 17.17 71.33 34.33 17.34
Qwen Reliability ($\uparrow$) 100.00 100.00 95.33 100.00 98.67 10.33 7.83 100.00
Generality ($\uparrow$) 99.94 81.81 86.69 95.31 98.64 7.00 6.44 99.34
Audio Locality ($\uparrow$) 67.40 90.49 83.36 83.29 68.00 87.47 82.82 70.04
Text Locality ($\uparrow$) 75.58 100.00 85.67 87.25 72.84 55.67 51.17 82.09
Portability ($\uparrow$) 24.00 49.92 26.75 26.25 26.58 28.50 27.58 27.84
AF Reliability ($\uparrow$) 100.00 100.00 100.00 99.92 99.92 25.00 58.83 100.00
Generality ($\uparrow$) 93.31 76.08 98.89 93.92 91.17 15.08 58.36 77.69
Audio Locality ($\uparrow$) 87.67 90.24 81.38 90.64 76.03 87.07 40.42 91.18
Text Locality ($\uparrow$) 97.00 100.00 95.83 98.67 81.17 81.17 76.67 96.08
Portability ($\uparrow$) 31.67 49.50 31.92 32.08 31.75 37.08 40.67 30.75
@@ 5.1 Single Editing
(a) Results for DeSTA.
(b) Results for Qwen.
(c) Results for AF.
Figure 3: Performances of editing methods on types of generality (left) and audio locality (right). Curves represent evaluation types.
@@
The main results of different editing methods on SAKE are shown in Table 2. Detailed results for each auditory attribute and relevant discussions are provided in Appendix F.
@@
Reliability: Parameter-updating methods are almost perfectly reliable, while IKE variants underperform. For all LALMs, parameter-updating methods (i.e., methods other than IKE variants) all achieve high reliability (mostly $\geq$95% and often $\approx$100%), showing their effectiveness in enforcing the edits. Among them, WISE consistently attains the best performance. In contrast, I-IKE and IE-IKE perform poorly. Interestingly, I-IKE outperforms IE-IKE on DeSTA and Qwen, despite the latter using additional auditory examples. We attribute this to those LALMs’ limited in-context learning ability: they struggle to handle multi-audio inputs and leverage examples, unlike in LLMs (Zheng et al., 2023) and LVLMs (Cheng et al., 2023; Huang et al., 2024) where such methods are effective.
Generality: Enforcing edits does not guarantee consistent generalization beyond the edited instances. Overall generality scores are consistently lower than reliability across all LALMs, indicating that successful edits on target instances do not readily extend to the equivalent neighborhoods. WISE and KE achieve the strongest average generality on DeSTA and AF, respectively, while FT (LLM) performs best on Qwen. Interestingly, FT (LLM) consistently outperforms FT (Audio) in average generality despite comparable reliability, suggesting that updating the LLM backbone yields better generalization than tuning the modality connector alone. We analyze the performances across types of generality in Sec. 5.2.
Audio Locality: Preserving unrelated auditory knowledge remains challenging for most editing methods. No method consistently dominates across LALMs. On average, audio locality scores of the methods are lower than reliability and generality, indicating that enforcing auditory edits often interferes with unrelated auditory knowledge. FT (LLM) shows lower audio locality than FT (Audio), suggesting that editing different model components induces different levels of collateral effects in the auditory domain. Together with the generality results, these observations point to a trade-off between generalizing edited knowledge and preserving unrelated auditory knowledge. Although IKE variants do not achieve high reliability or generality, they preserve audio locality to a certain extent, likely because limited in-context learning capability constrains the propagation of the edits, albeit at the cost of weaker editing effectiveness. We discuss different types of audio locality in Sec. 5.2.
Text locality: Largely preserved when edits avoid modifying the LLM backbone. FT (Audio) maintains perfect text locality, as it only updates the modality connector and leaves text-only behavior unaffected. In contrast, FT (LLM) substantially degrades text locality on most LALMs due to direct backbone updates. KE and MEND preserve text locality reasonably well through regularization. Other methods show larger drops, indicating that safeguarding textual capabilities requires explicit considerations during editing.
Table 3: Standard deviation (Std.) and range of portability scores across knowledge categories for the top-two methods per model.
Model Method Std. (%) Range (%)
DeSTA I-IKE 12.66 45.13
FT (Audio) 22.64 88.18
Qwen FT (Audio) 17.09 62.27
I-IKE 15.45 72.37
AF FT (Audio) 16.64 65.46
IE-IKE 14.51 54.40
Portability: Propagating auditory edits to connected knowledge remains an open challenge. Overall, current editing methods do not guarantee portability when modifying auditory attribute knowledge. While no approach consistently excels, FT (Audio) exhibits comparatively more balanced portability behavior across settings. This observation suggests that edits applied through the modality connector can interact differently with related knowledge than backbone updates, though portability remains far from resolved. More broadly, the limited portability observed across methods highlights a practical challenge for scenarios where modifying an auditory attribute is expected to induce coherent updates to a broader set of related knowledge.
We further analyze portability by categorizing the related knowledge in the portability set. For each LALM, we select the two best-performing editing methods and report the standard deviation and range of their portability scores across knowledge categories in Table 3, with full results provided in Appendix I. The large variation across categories shows that portability failures are not uniform: current editing methods can successfully update certain types of related knowledge while leaving others unchanged, indicating that portability exhibits substantial variation across knowledge categories rather than an all-or-nothing outcome.
Summary. While most methods can enforce the desired behavior on edited instances, they struggle to generalize to equivalent ones, propagate edits to connected knowledge, and preserve unrelated knowledge within the same attribute, particularly in the auditory domain. Fine-tuning remains a strong baseline, with FT (Audio) showing consistently competitive performance. These results underscore that editing auditory attribute knowledge remains highly challenging and calls for auditory-specific editing methods.
(a) Results for DeSTA.
(b) Results for Qwen.
(c) Results for AF.
@@
5.2 Detailed Analyses of Different Types of Generality and Audio Locality
We compare editing performance across different types of generality and audio locality in Figure 3. Unlike prior works in vision domains (Cheng et al., 2023; Zhang et al., 2025; Huang et al., 2024; Du. et al., 2025) that do not explicitly differentiate multiple forms of knowledge generalization and locality, our design decomposes both dimensions into distinct types, enabling more fine-grained analyses of how edits generalize and how unrelated auditory knowledge is preserved.
For generality, editing methods generally perform well on type 1 (textual neighborhood) but degrade on type 2, which requires generalization to acoustically similar yet non-identical inputs with the same labels. Type 3, which combines both textual and auditory neighborhoods, is consistently the most challenging, indicating that extending edited knowledge along the auditory dimension remains substantially more difficult than along the textual one.
For audio locality, type 2 probes knowledge within the same attribute but unrelated to the edit, and proves hardest to preserve compared to knowledge of other attributes (type 1) or the edited target itself (type 3). This pattern shows that modifying one aspect of an auditory attribute can inadvertently affect other aspects of the same attribute, revealing a form of intra-attribute entanglement in auditory attribute knowledge. Consistent with this observation, FT (LLM) exhibits substantially lower locality on type 2 than FT (Audio), despite its stronger generalization performance, suggesting a tension between generalizing edits and preserving unrelated knowledge within the same attribute. We provide a more detailed discussion of this phenomenon in Appendix H.
Type 4 assesses the preservation of general auditory processing ability. Most methods show clear degradation, suggesting that targeted edits often interfere with broader auditory competence. In contrast, KE and MEND better preserve original behavior, likely due to their explicit locality-oriented regularization during hypernetwork training.
@@ 5.3 Sequential Editing
Figure 4 shows the results on sequential editing.
Significant forgetting under sequential editing in reliability and generality. Most methods show clear declines in reliability and generality as the edit gap increases, indicating substantial forgetting of previously edited auditory knowledge. In contrast, MEND and KE degrade rapidly after only a few edits in most cases, except for MEND on AF. We attribute this behavior to their sensitivity to sequential updates, which we will discuss later.
Sequential editing can induce degeneration. Some editing methods exhibit degeneration, characterized by repetitive or nonsensical outputs (Holtzman et al., 2020; Gupta et al., 2024; Yang et al., 2024b; Hsueh et al., 2024). This phenomenon is widely observed in many cases, e.g., MEND on DeSTA and Qwen, KE on all models, and UnKE on AF. These results suggest that such methods are particularly sensitive to sequential editing, where successive updates interfere with one another and induce model collapse. This highlights a key limitation of current editing approaches. Statistics and qualitative examples of degeneration are provided in Appendix J. In contrast, methods like FT (Audio) show substantially less degeneration after sequential updates, showing greater robustness.
Strength-stability trade-off in IKE variants. Although weaker in single-edit settings, the IKE variants remain relatively stable under larger gaps, with markedly smaller degradation than parameter-updating methods; in particular, I-IKE achieves the strongest long-term reliability and generality on DeSTA.
No significant locality degradation under sequential editing for most methods. Most methods do not exhibit the sharp performance drops observed in reliability and generality, indicating that sequential editing has a comparatively weaker impact on locality. Nevertheless, as discussed in Sec. 5.2, substantial room for improvement remains, since none of the methods consistently preserve all kinds of unrelated knowledge. Notably, KE shows a pronounced decline at larger edit gaps on all LALMs due to degeneration issues.
Portability consistently falls below random baseline under sequential editing for most methods. Under sequential editing, most methods demonstrate persistently poor portability, consistent with the findings reported in Sec. 5.1. In particular, most methods perform below the random baseline (27.5%) even at small edit gaps, indicating that their ability to transfer edits to connected knowledge is limited. As a result, increasing the edit gap does not lead to substantial additional degradation or noticeable fluctuations, since portability performance is already near its lower bound.
FT (Audio) remains a strong baseline under sequential editing. Among all methods, FT (Audio) demonstrates relatively strong robustness across most LALMs and metrics, with notable degradation observed only in reliability and generality on DeSTA. Combined with the results in Sec. 5.1, where FT (Audio) achieves the most balanced performance in single-edit settings, these findings indicate that FT (Audio) remains a competitive baseline for auditory knowledge editing, despite its conceptual simplicity.